b only
v(t) Set up the differential equation and solve for the voltage in the circuit shown to the right for the following component values. a) i(t)-3x(t), b)i(t)-Su(t), R-200 Ω, L-4nH, i1(0)-0A R-1000 Ω, L-40 ml, iL(0)-2A
Answer #1

The given circuit will be easily solved in the Laplace domain.

In the Laplace domain, the circuit will be represented as

From the circuit, we have

r(t) = Ldi (t)

Applying Laplace transform on the above equation,

V(s)=sLI_L(s)-i_L(0) ....... Eq.1

According to the voltage-division principle,

l(s) = R+sL1(s) ...... Eq.2

From Eq.1 and Eq.2,

V(s)=sLrac{R}{R+sL}I(s)-i_L(0)

Rightarrow V(s)=rac{sLR}{R+sL}I(s)-i_L(0) ........ Eq.3

Rightarrow (R+sL)V(s)=I(s)(R+sL)-i_L(0)(R+sL)

Applying inverse Laplace transforms

du(t) di(t).di(t)) dt

Ldv(t) di(t) L di(t)) R dt R dt2

(a) i(t)=3u(t); R=200Omega; L=4mH; i_L(0)=0A

1 (s) =

From Eq.3

Rightarrow V(s)=rac{s4 imes10^{-3} imes200}{200+s4 imes10^{-3}}rac{3}{s}-0

(s)-(2000 0004s) 3

Rightarrow V(s)=rac{2.4}{(200+0.004s)}

600 V(s) = (s + 50000)

L^{-1}left[rac{1}{s+a} ight ]=e^{-at}

Applying inverse Laplace transform

v(t)=600e^{-50000t}V

(b) i(t)=5u(t); R=1000Omega; L=40mH; i_L(0)=2A

I(s)-

From Eq.3

Rightarrow V(s)=rac{s imes40 imes10^{-3} imes1000}{1000+s40 imes10^{-3}}rac{5}{s}-2

200 1000 ± 0.04s V(s)-

Rightarrow V(s)=rac{5000}{25000+s}-2

Applying inverse Laplace transforms

v(t)=5000e^{-25000t}-2delta(t)